# QVGen: Pushing the Limit of Quantized Video Generative Models

### v3

@@ Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., $4$-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization $\bm{\gamma}$} to identify and decay low-contributing components. This strategy retains performance while zeroing out inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3\text{B}\sim 14\text{B}$, show that QVGen is the first to reach full-precision comparable quality under $4$-bit settings. Moreover, it significantly outperforms existing methods. For instance, our $3$-bit CogVideoX-22B achieves improvements of +25.28+25.28 in Dynamic Degree and +8.43+8.43 in Scene Consistency on VBench.
@@ 3.1 Improving Convergence with Auxiliary Modules $\Phi$
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 3.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 3.2): (i) SVD to identify the low-impact components in $\Phi$; (ii) A rank-based regularization $\bm{\gamma}$} to decay the identified components to $\varnothing$. Overall detailed algorithm can be found in Sec. B
@@
A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes (Takase et al., 2024; Xie et al., 2024). Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes) (Kumar et al., 2025; Li et al., 2024b). Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This module is trained to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
@@
where $\Phi({\mathcal{Q}}_{b}({\mathbf{X}}))={\mathbf{W}}_{\Phi}{\mathcal{Q}}_{b}({\mathbf{X}})$. Here, ${\mathbf{W}}_{\Phi}$ is initialized before QAT by the weight quantization error, defined as ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$.
@@ 3.2 Progressively Shrinking $\Phi$ via Rank-Decay
$\begin{array}[]{ll}{\mathbf{W}}_{\Phi}=\sum_{s=1}^{d}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{T},\end{array}$ (10)
@@
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small (Zhang et al., 2015; Yang et al., 2020). Hence, as training proceeds, only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{T}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
@@
where ${\mathbf{L}}=[\sqrt{\sigma_{1}}{\bm{u}}_{1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]\in\mathbb{R}^{n\times r}$ and ${\mathbf{R}}=[\sqrt{\sigma_{1}}{\bm{v}}_{1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{T}\in\mathbb{R}^{r\times m}$ for a given rank $r$. In practice, we set $r\ll d$ to reduce training costs, as ${\mathbf{W}}_{\Phi}$ already exhibits a non-negligible number of small singular values before QAT. Consequently, $\sqrt{\sigma_{s}}{\bm{u}}_{s}$ and $\sqrt{\sigma_{s}}{\bm{v}}_{s}$ are the components in ${\mathbf{W}}_{\Phi}$ represent the $s$-th level of contribution. Then, with a rank-based regularization $\bm{\gamma}$} applied, the forward computation of a quantized linear layer during training is modified as:
@@
Here, $u$ follows a cosine annealing schedule that decays from $1$ to 0, $\lambda\in(0,1]$ represents the shrinking ratio, and $\odot$ denotes element-wise multiplication. Eq. (12) and Eq. (13) allow us to progressively eliminate the low-impact components of $\Phi$ (i.e., $\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{u}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]$ and $\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{v}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{T}$). Once $u$ reaches 0, we truncate $\{{\mathbf{L}},{\mathbf{R}}\}$ to $\{{\mathbf{L}}^{\prime},{\mathbf{R}}^{\prime}\}$, and rewrite ${\mathbf{W}}_{\Phi}$ as:
@@ 4.2 Performance Analysis
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-55B and Wan 1414B, at $720p$ resolution. As illustrated in Tab. 2, our $3$-bit and $4$-bit models follow the same pattern seen in smaller models (Tab. 1). However, $3$-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency,
(a) CogVideoX1.5-55(Yang et al., 2025)
(b) Wan 1414(WanTeam et al., 2025)
underscoring the challenge of pushing these larger models to $3$ bits. In Fig. 5, we further assess the models with VBench-2.0; the W4A4 DMs incur only negligible overall performance loss.
@@ 4.4 Efficiency Discussion
Although previous research (Zhang et al., 2025b) mentioned that the current $3D$ full-attention video occupies significant computation in video generation, our QVGen is orthogonal to works that focus on accelerating $3D$ attention and can deliver notable speedups to models that already employ those techniques. For example, when SVG (Xi et al., 2025) is paired with W4A4 QVGen, the model runs $1.70\times$ and $2.63\times$ faster, compared with $1.31\times$ and $1.68\times$ for SVG alone, as shown in Fig. 6 (c).
@@ Appendix H Comparison with baselines on Additional Metrics
Table L: Performance comparison across different methods on VBench (Huang et al., 2024b). We employ W4A4 CogVideo-X 55(Yang et al., 2025).

### v4

@@ Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., $4$-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization $\bm{\gamma}$} to identify and decay low-contributing components. This strategy retains performance while zeroing out additional inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3\text{B}\sim 14\text{B}$, show that QVGen is the first to reach full-precision comparable quality under $4$-bit settings. Moreover, it significantly outperforms existing methods. For instance, our $3$-bit CogVideoX-22B achieves improvements of +25.28+25.28 in Dynamic Degree and +8.43+8.43 in Scene Consistency on VBench.
@@ 3 QVGen
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 3.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 3.2): (i) SVD to identify the low-impact components in $\Phi$; (ii) A rank-based regularization $\bm{\gamma}$} to decay the identified components to $\varnothing$. A detailed procedure can be found in Sec. B.
@@ 3.1 Improving Convergence with Auxiliary Modules $\Phi$
A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes (Takase et al., 2024; Xie et al., 2024). Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes(Kumar et al., 2025; Li et al., 2024b). Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This trainable module aims to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
@@
where $\Phi({\mathcal{Q}}_{b}({\mathbf{X}}))={\mathbf{W}}_{\Phi}{\mathcal{Q}}_{b}({\mathbf{X}})$. Here, ${\mathbf{W}}_{\Phi}$ is initialized before QAT by the weight quantization error, defined as ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$. More initialization approaches for ${\mathbf{W}}_{\Phi}$ can be found in Sec. J.2.
@@ 3.2 Progressively Shrinking $\Phi$ via Rank-Decay
$\begin{array}[]{ll}{\mathbf{W}}_{\Phi}=\sum_{s=1}^{d}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{\top},\end{array}$ (10)
@@
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small (Zhang et al., 2015; Yang et al., 2020). Hence, as training proceeds, only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{\top}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
@@
where ${\mathbf{L}}=[\sqrt{\sigma_{1}}{\bm{u}}_{1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]\in\mathbb{R}^{n\times r}$ and ${\mathbf{R}}=[\sqrt{\sigma_{1}}{\bm{v}}_{1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{\top}\in\mathbb{R}^{r\times m}$ for a given rank $r$. In practice, we set $r\ll d$ to reduce training costs, as ${\mathbf{W}}_{\Phi}$ already exhibits a non-negligible number of small singular values before QAT. Consequently, $\sqrt{\sigma_{s}}{\bm{u}}_{s}$ and $\sqrt{\sigma_{s}}{\bm{v}}_{s}$ are the components in ${\mathbf{W}}_{\Phi}$ represent the $s$-th level of contribution. Then, with a rank-based regularization $\bm{\gamma}$} applied, the forward computation of a quantized linear layer during training is modified as:
@@
Here, $u$ follows a cosine annealing schedule that decays from $1$ to 0, $\lambda\in(0,1]$ represents the shrinking ratio, and $\odot$ denotes element-wise multiplication. Eq. (12) and Eq. (13) allow us to progressively eliminate the low-impact components of $\Phi$ (i.e., $\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{u}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]$ and $\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{v}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{\top}$). Once $u$ reaches 0, we truncate $\{{\mathbf{L}},{\mathbf{R}}\}$ to $\{{\mathbf{L}}^{\prime},{\mathbf{R}}^{\prime}\}$, and rewrite ${\mathbf{W}}_{\Phi}$ as:
@@ 4.2 Performance Analysis
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-55B and Wan 1414B, at $720p$ resolution. As illustrated in Tab. 2, our $3$-bit and $4$-bit models follow the same pattern seen in smaller models (Tab. 1). However, $3$-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency, underscoring the challenge of pushing these larger models to $3$ bits. In Fig. 5, we further assess the models with VBench-2.0; the W4A4 DMs incur only negligible overall performance loss.
@@ 4.4 Efficiency Discussion
Although previous research (Zhang et al., 2025b) mentioned that the current $3D$ full-attention occupies significant computation in video generation, our QVGen is orthogonal to works that focus on accelerating $3D$ attention and can deliver notable speedups to models that already employ those techniques. For example, when SVG (Xi et al., 2025) is paired with W4A4 QVGen, the model runs $1.70\times$ and $2.63\times$ faster, compared with $1.31\times$ and $1.68\times$ for SVG alone, as shown in Fig. 6 (c).
@@ Appendix H Comparison with baselines on Additional Metrics
Table L: Performance comparison for huge DMs across different methods on VBench (Huang et al., 2024b). We employ W4A4 CogVideo-X 55(Yang et al., 2025) here.
@@ Appendix I Comparison with Baselines for Huge DMs
Appendix M The Use of Large Language Models
We acknowledge the use of large language models (LLMs), such as OpenAI’s GPT-5, as a writing-assistance tool in this work. Their role was strictly limited to proofreading and rephrasing sentences to enhance linguistic quality, without any contribution to the research ideation or experimental results.